In a triangle ABC,with angels A,B and C and sides AB,BC,AC, angle B is a right (90)angle. If the sin of angle A is 0.5 and side BC is 8 inches long, what is the length of side AC?
Answer by rothauserc(4718)

since we have a right triangle, we know
:
sin(A) = BC / AC
:
we know sin(A) = (1/2) and BC = 8 inches, therefore
:
(1 / 2) = (8 / AC)
:
cross multiply the two fractions since they are equal
:
AC = 16 inches
: